\begin{thebibliography}{10}

\bibitem{ACHLIOPTAS}
Dimitris Achlioptas.
\newblock Database-friendly random projections: {J}ohnson-{L}indenstrauss with
  binary coins.
\newblock volume~66, pages 671--687. 2003.
\newblock Special issue on PODS 2001 (Santa Barbara, CA).

\bibitem{AILONCHAZELLE}
Nir Ailon and Bernard Chazelle.
\newblock Approximate nearest neighbors and the fast {J}ohnson-{L}indenstrauss
  transform.
\newblock In {\em S{TOC}'06: {P}roceedings of the 38th {A}nnual {ACM}
  {S}ymposium on {T}heory of {C}omputing}, pages 557--563. ACM, New York, 2006.

\bibitem{AILONCHAZELLE2009}
Nir Ailon and Bernard Chazelle.
\newblock The fast {J}ohnson-{L}indenstrauss transform and approximate nearest
  neighbors.
\newblock {\em SIAM J. Comput.}, 39(1):302--322, 2009.

\bibitem{AILONLIBERTY}
Nir Ailon and Edo Liberty.
\newblock Fast dimension reduction using {R}ademacher series on dual {BCH}
  codes.
\newblock {\em Discrete Comput. Geom.}, 42(4):615--630, 2009.

\bibitem{AILONLIBERTY2}
Nir Ailon and Edo Liberty.
\newblock An almost optimal unrestricted fast {J}ohnson-{L}indenstrauss
  transform.
\newblock {\em ACM Trans. Algorithms}, 9(3):12 Pages, 2013.

\bibitem{ALON}
Noga Alon.
\newblock Problems and results in extremal combinatorics. {I}.
\newblock {\em Discrete Math.}, 273(1-3):31--53, 2003.

\bibitem{ALONKLARTAG}
Noga Alon and Bo'az Klartag.
\newblock Optimal compression of approximate inner products and dimension
  reduction.
\newblock In {\em 58th {A}nnual {IEEE} {S}ymposium on {F}oundations of
  {C}omputer {S}cience---{FOCS} 2017}, pages 639--650. IEEE Computer Soc., Los
  Alamitos, CA, 2017.

\bibitem{ARRIAGAVEMPALA}
Rosa~I. Arriaga and Santosh Vempala.
\newblock An algorithmic theory of learning: robust concepts and random
  projection.
\newblock In {\em 40th {A}nnual {S}ymposium on {F}oundations of {C}omputer
  {S}cience ({N}ew {Y}ork, 1999)}, pages 616--623. IEEE Computer Soc., Los
  Alamitos, CA, 1999.

\bibitem{BARTALRECHTSCHULMAN}
Yair Bartal, Ben Recht, and Leonard~J. Schulman.
\newblock Dimensionality reduction: beyond the {J}ohnson-{L}indenstrauss bound.
\newblock In {\em Proceedings of the {T}wenty-{S}econd {A}nnual {ACM}-{SIAM}
  {S}ymposium on {D}iscrete {A}lgorithms}, pages 868--887. SIAM, Philadelphia,
  PA, 2011.

\bibitem{BECCHETTIBURYCOHENADDADGRANDONISCHWIEGELSHOHN}
Luca Becchetti, Marc Bury, Vincent Cohen-Addad, Fabrizio Grandoni, and Chris
  Schwiegelshohn.
\newblock Oblivious dimension reduction for {$k$}-means: beyond subspaces and
  the {J}ohnson-{L}indenstrauss lemma.
\newblock In {\em S{TOC}'19---{P}roceedings of the 51st {A}nnual {ACM} {SIGACT}
  {S}ymposium on {T}heory of {C}omputing}, pages 1039--1050. ACM, New York,
  2019.

\bibitem{BOURGAIN}
J.~Bourgain.
\newblock On {L}ipschitz embedding of finite metric spaces in {H}ilbert space.
\newblock {\em Israel J. Math}, 52(1-2):46--52, 1985.

\bibitem{BOURGAINTZAFRIRI}
J.~Bourgain and L.~Tzafriri.
\newblock Invertibility of ``large'' submatrices with applications to the
  geometry of {B}anach spaces and harmonic analysis.
\newblock {\em Israel J. Math.}, 57(2):137--224, 1987.

\bibitem{BOURGAINTZAFRIRI2}
J.~Bourgain and L.~Tzafriri.
\newblock Restricted invertibility of matrices and applications.
\newblock In {\em Analysis at Urbana, Vol. II (Urbana, IL, 1986–1987), volume
  138 of London Math. Soc. Lecture Note Ser.}, pages 67--107. Cambridge Univ.
  Press, Cambridge, 1989.

\bibitem{CASAZZA}
Peter~G. Casazza.
\newblock The simplified version of the {S}pielman and {S}rivastava algorithm
  for proving the {B}ourgain-{T}zafriri restricted invertiblity theorem.
\newblock {\em arXiv:1208.4013v2, 21 August}, 2012.

\bibitem{CASAZZAPFANDER}
Peter~G. Casazza and G\"{o}tz~E. Pfander.
\newblock Infinite dimensional restricted invertibility.
\newblock {\em J. Funct. Anal.}, 263(12):3784--3803, 2012.

\bibitem{CASAZZATREMAIN}
Peter~G. Casazza and Janet~C. Tremain.
\newblock Revisiting the {B}ourgain-{T}zafriri restricted invertibility
  theorem.
\newblock {\em Oper. Matrices}, 3(1):97--110, 2009.

\bibitem{CHERVOVFALQUI}
A.~Chervov and G~Falqui.
\newblock {M}anin matrices and and {T}alalaev's formula.
\newblock {\em J. Phys. A: Math. Theor.}, 41:1--28, 2008.

\bibitem{CHERVOVFALQUIRUBTSOV}
A.~Chervov, G.~Falqui, and V.~Rubtsov.
\newblock Algebraic properties of {M}anin matrices - 1.
\newblock {\em Advances in Applied Mathematics}, 43:239--315, 2009.

\bibitem{CHRETIENDARSES}
St\'{e}phane Chretien and Sebastien Darses.
\newblock On a new method for controlling the entire spectrum in the problem of
  column subset selection.
\newblock {\em Expo. Math.}, 37(3):314--321, 2019.

\bibitem{COHENELDERMUSCOMUSCOPERSU}
Michael~B. Cohen, Sam Elder, Cameron Musco, Christopher Musco, and Madalina
  Persu.
\newblock Dimensionality reduction for k-means clustering and low rank
  approximation.
\newblock In {\em S{TOC}'15---{P}roceedings of the 2015 {ACM} {S}ymposium on
  {T}heory of {C}omputing}, pages 163--172. ACM, New York, 2015.

\bibitem{DASGUPTAKUMARSARLOS}
Anirban Dasgupta, Ravi Kumar, and Tamas Sarlos.
\newblock A sparse {J}ohnson-{L}indenstrauss transform.
\newblock In {\em S{TOC}'10---{P}roceedings of the 2010 {ACM} {I}nternational
  {S}ymposium on {T}heory of {C}omputing}, pages 341--350. ACM, New York, 2010.

\bibitem{DASGUPTAGUPTA}
Sanjoy Dasgupta and Anupam Gupta.
\newblock An elementary proof of a theorem of {J}ohnson and {L}indenstrauss.
\newblock {\em Random Structures Algorithms}, 22(1):60--65, 2003.

\bibitem{DEDEZALAURENT}
Michel~Marie Deza and Monique Laurent.
\newblock {\em Geometry of Cuts and Metrics}, volume~1 of {\em Algorithms and
  Combinatorics}.
\newblock Springer, Berlin, 1997.

\bibitem{FOUCARTRAUHUT}
Simon Foucart and Holger Rauhut.
\newblock {\em A mathematical introduction to compressive sensing}.
\newblock Applied and Numerical Harmonic Analysis. Birkh\"{a}user/Springer, New
  York, 2013.

\bibitem{FRANKLMAEHARA}
P.~Frankl and H.~Maehara.
\newblock The {J}ohnson-{L}indenstrauss lemma and the sphericity of some
  graphs.
\newblock {\em J. Combin. Theory Ser. B}, 44(3):355--362, 1988.

\bibitem{FREKSEN}
Casper~Benjamin Freksen.
\newblock An introduction to {J}ohnson-{L}indenstrauss transforms.
\newblock {\em arXiv:2103.00564v1 [cs.DS] 28 February}, 2021.

\bibitem{GIPSON}
Philip~M. Gipson.
\newblock Invariant basis number for {C}*-algebras.
\newblock {\em Illinois Journal of Mathematics}, 59(1):85--98, 2015.

\bibitem{GLUSKINOLEVSKII}
E.~Gluskin and A.~Olevskii.
\newblock Invertibility of sub-matrices and the octahedron width theorem.
\newblock {\em Israel J. Math.}, 186:61--68, 2011.

\bibitem{INDYKNAOR}
P.~Indyk and A.~Naor.
\newblock Nearest-neighbor-preserving embeddings.
\newblock {\em ACM Transactions on Algorithms}, 3(3):1--12, 2007.

\bibitem{INDYK2000}
Piotr Indyk.
\newblock Dimensionality reduction techniques for proximity problems.
\newblock In {\em Proceedings of the {E}leventh {A}nnual {ACM}-{SIAM}
  {S}ymposium on {D}iscrete {A}lgorithms ({S}an {F}rancisco, {CA}, 2000)},
  pages 371--378. ACM, New York, 2000.

\bibitem{INDYKMOTWANI}
Piotr Indyk and Rajeev Motwani.
\newblock Approximate nearest neighbors: towards removing the curse of
  dimensionality.
\newblock In {\em S{TOC} '98 ({D}allas, {TX})}, pages 604--613. ACM, New York,
  1999.

\bibitem{JACQUES}
Laurent Jacques.
\newblock A quantized {J}ohnson-{L}indenstrauss lemma: the finding of
  {B}uffon's needle.
\newblock {\em IEEE Trans. Inform. Theory}, 61(9):5012--5027, 2015.

\bibitem{JAYRAMWOODRUFF}
T.~S. Jayram and D.~P. Woodruff.
\newblock Optimal bounds for {J}ohnson-{L}indenstrauss transforms and streaming
  problems with subconstant error.
\newblock {\em ACM Transactions on Algorithms}, 9(3):1--17, 2013.

\bibitem{JOHNSONLINDENSTRAUSS}
William~B. Johnson and Joram Lindenstrauss.
\newblock Extensions of {L}ipschitz mappings into a {H}ilbert space.
\newblock In {\em Conference in modern analysis and probability ({N}ew {H}aven,
  {C}onn., 1982)}, volume~26 of {\em Contemp. Math.}, pages 189--206. Amer.
  Math. Soc., Providence, RI, 1984.

\bibitem{KANEMEKANELSON}
Daniel Kane, Raghu Meka, and Jelani Nelson.
\newblock Almost optimal explicit {J}ohnson-{L}indenstrauss families.
\newblock In {\em Approximation, randomization, and combinatorial
  optimization}, volume 6845 of {\em Lecture Notes in Comput. Sci.}, pages
  628--639. Springer, Heidelberg, 2011.

\bibitem{KANENELSON}
Daniel~M. Kane and Jelani Nelson.
\newblock Sparser {J}ohnson-{L}indenstrauss transforms.
\newblock {\em J. ACM}, 61(1):Art. 4, 23, 2014.

\bibitem{KAPLANSKY}
Irving Kaplansky.
\newblock Modules over operator algebras.
\newblock {\em Amer. J. Math.}, 75:839--858, 1953.

\bibitem{KLARTAGMENDELSON}
B.~Klartag and S.~Mendelson.
\newblock Empirical processes and random projections.
\newblock {\em J. Funct. Anal.}, 225(1):229--245, 2005.

\bibitem{KRAHMERWARD}
Felix Krahmer and Rachel Ward.
\newblock New and improved {J}ohnson-{L}indenstrauss embeddings via the
  restricted isometry property.
\newblock {\em SIAM J. Math. Anal.}, 43(3):1269--1281, 2011.

\bibitem{MAHESHKRISHNA}
K.~Mahesh Krishna.
\newblock Modular {W}elch bounds with applications.
\newblock {\em arXiv:2201.00319 [math.OA] 2 January}, 2022.

\bibitem{LARSENNELSON}
Kasper~Green Larsen and Jelani Nelson.
\newblock The {J}ohnson-{L}indenstrauss lemma is optimal for linear
  dimensionality reduction.
\newblock In {\em 43rd {I}nternational {C}olloquium on {A}utomata, {L}anguages,
  and {P}rogramming}, volume~55 of {\em LIPIcs. Leibniz Int. Proc. Inform.},
  pages Art. No. 82, 11. Schloss Dagstuhl. Leibniz-Zent. Inform., Wadern, 2016.

\bibitem{LARSENNELSON2017}
Kasper~Green Larsen and Jelani Nelson.
\newblock Optimality of the {J}ohnson-{L}indenstrauss lemma.
\newblock In {\em 58th {A}nnual {IEEE} {S}ymposium on {F}oundations of
  {C}omputer {S}cience---{FOCS} 2017}, pages 633--638. IEEE Computer Soc., Los
  Alamitos, CA, 2017.

\bibitem{LIBERTYAILONSINGER}
Edo Liberty, Nir Ailon, and Amit Singer.
\newblock Dense fast random projections and lean {W}alsh transforms.
\newblock {\em Discrete Comput. Geom.}, 45(1):34--44, 2011.

\bibitem{LINIALLONDONRABINOVICH}
N.~Linial, E.~London, and Y.~Rabinovich.
\newblock The geometry of graphs and some of its algorithmic applications.
\newblock {\em Combinatorica}, 15(2):215--245, 1995.

\bibitem{MARCUSSPIELMANSRIVASTAVA}
Adam~W. Marcus, Daniel~A. Spielman, and Nikhil Srivastava.
\newblock Interlacing families {III}: {S}harper restricted invertibility
  estimates.
\newblock {\em Israel J. Math.}, 247(2):519--546, 2022.

\bibitem{MATOUSEKBOOK}
Jiri Matousek.
\newblock {\em Lectures on discrete geometry}, volume 212 of {\em Graduate
  Texts in Mathematics}.
\newblock Springer-Verlag, New York, 2002.

\bibitem{NAOR}
Assaf Naor.
\newblock Sparse quadratic forms and their geometric applications [following
  {B}atson, {S}pielman and {S}rivastava].
\newblock {\em Asterisque}, 348:189--217, 2012.

\bibitem{NAORYOUSSEF}
Assaf Naor and Pierre Youssef.
\newblock Restricted invertibility revisited.
\newblock In {\em A journey through discrete mathematics}, pages 657--691.
  Springer, Cham, 2017.

\bibitem{OSTROVSKII}
Mikhail~I. Ostrovskii.
\newblock {\em Metric Embeddings: {B}ilipschitz and Coarse Embeddings into
  {B}anach Spaces}, volume~49 of {\em De Gruyter Studies in Mathematics}.
\newblock De Gruyter, Berlin, 2013.

\bibitem{PASCHKE}
William~L. Paschke.
\newblock Inner product modules over {$B^{\ast} $}-algebras.
\newblock {\em Trans. Amer. Math. Soc.}, 182:443--468, 1973.

\bibitem{RAVICHANDRAN}
Mohan Ravichandran.
\newblock Principal submatrices, restricted invertibility, and a quantitative
  {G}auss–{L}ucas theorem.
\newblock {\em International Mathematics Research Notices}, (15):4809--4832,
  2018.

\bibitem{RIEFFEL}
Marc~A. Rieffel.
\newblock Induced representations of {$C^{\ast} $}-algebras.
\newblock {\em Advances in Math.}, 13:176--257, 1974.

\bibitem{SPIELMANSRIVASTAVA}
Daniel~A. Spielman and Nikhil Srivastava.
\newblock An elementary proof of the restricted invertibility theorem.
\newblock {\em Israel J. Math.}, 190:83--91, 2012.

\bibitem{SRIVASTAVA}
Nikhil Srivastava.
\newblock Spectral sparsification and restricted invertibility.
\newblock {\em Ph.D. Thesis, Yale University}, pages (vii)+70, 2010.

\bibitem{TROPP2}
Joel~A. Tropp.
\newblock The random paving property for uniformly bounded matrices.
\newblock {\em Studia Math.}, 185(1):67--82, 2008.

\bibitem{TROPP}
Joel~A. Tropp.
\newblock Column subset selection, matrix factorization, and eigenvalue
  optimization.
\newblock In {\em Proceedings of the {T}wentieth {A}nnual {ACM}-{SIAM}
  {S}ymposium on {D}iscrete {A}lgorithms}, pages 978--986. SIAM, Philadelphia,
  PA, 2009.

\bibitem{VENKATASUBRAMANIANWANG}
Suresh Venkatasubramanian and Qiushi Wang.
\newblock The {J}ohnson-{L}indenstrauss transform: An empirical study.
\newblock In {\em Proceedings of the Meeting on Algorithm Engineering and
  Expermiments}, pages 164--173. SIAM, San Francisco, California, 2011.

\bibitem{VERSHYNIN}
R.~Vershynin.
\newblock John's decompositions: selecting a large part.
\newblock {\em Israel J. Math.}, 122:253--277, 2001.

\bibitem{YOUSSEF}
Pierre Youssef.
\newblock A note on column subset selection.
\newblock {\em International Mathematics Research Notices}, (23):6431--6447,
  2013.

\bibitem{YOUSSEF2}
Pierre Youssef.
\newblock Restricted invertibility and the {B}anach-{M}azur distance to the
  cube.
\newblock {\em Mathematika}, 60(1):201--218, 2014.

\end{thebibliography}

\documentclass{amsart}
\usepackage{setspace}
\usepackage{a4}
\usepackage{amssymb,amsmath,amsthm,latexsym}
\usepackage{amsfonts}
\usepackage{amsfonts}
\usepackage{graphicx}
\usepackage{textcomp}
\usepackage{cite}
\usepackage{enumerate}
\usepackage[mathscr]{euscript}
\usepackage{mathtools}
\newtheorem{theorem}{Theorem}[section]

\newtheorem{conjecture}[theorem]{Conjecture}

\newtheorem{problem}[theorem]{Problem}

\newtheorem{remark}[theorem]{Remark}

\title{This is the title}
\usepackage{hyperref}
\usepackage{enumerate}
\usepackage{mathtools}
\usepackage{amsmath}
\usepackage{tikz}
\usepackage{hyperref}
\usepackage{graphicx}
\usepackage{textcomp}
\usetikzlibrary{cd}
\usepackage{fancyhdr}

\begin{document}
\begin{center}
{\bf{MODULAR BOURGAIN-TZAFRIRI RESTRICTED INVERTIBILITY CONJECTURES AND JOHNSON-LINDENSTRAUSS FLATTENING   CONJECTURE}}\\
\vspace{0.2cm}
\hrule 
\vspace{0.2cm}
\textbf{K. MAHESH KRISHNA} \\
Post Doctoral Fellow \\
Statistics and Mathematics Unit\\
Indian Statistical Institute, Bangalore Centre\\
Karnataka 560 059 India\\
Email: kmaheshak@gmail.com \\
Date: \today
\end{center}

\hrule
\vspace{0.5cm}
%--------------------------------------
\textbf{Abstract}:  We recently formulated  important Modular Bourgain-Tzafriri Restricted Invertibility Conjectures and Modular Johnson-Lindenstrauss Flattening  Conjecture in the Appendix of  \textit{[arXiv: 2207.12799.v1]}. For the sake of wide accessibility we give a self-contained treatment of them.

\textbf{Keywords}: C*-algebra, Bourgain-Tzafriri Restricted Invertibility Theorem, Manin Matrix, Johnson-Lindenstrauss Flattening Lemma.

\textbf{Mathematics Subject Classification (2020)}: 46L05, 15A60.
\hrule
\tableofcontents
\hrule

\section{Modular Bourgain-Tzafriri Restricted Invertibility Conjectures}
 Let $d\in \mathbb{N}$, $ \mathbb{K}=\mathbb{C}$ or $\mathbb{R}$ and $\mathbb{K}^d$ be the standard $d$-dimensional Hilbert space with canonical orthonormal basis $\{e_j\}_{j=1}^d$.  Following result obtained by  Prof. Bourgain and Prof. Tzafriri  in 1987 is one of the highly useful  result of \textbf{20 century Mathematics} which combines algebraic property of a matrix  (invertibility) with the analytic property (norm) \cite{BOURGAINTZAFRIRI}.
\begin{theorem}\cite{BOURGAINTZAFRIRI, BOURGAINTZAFRIRI2}\label{BOURGAINTZARIRITHEOREM}  \textbf{(Bourgain-Tzafriri Restricted Invertibility Theorem)
		There are universal constants $A>0$, $c>0$ satisfying the following property. If  $d \in \mathbb{N}$, and  $T:\mathbb{K}^d \to \mathbb{K}^d$ is a linear operator   with  $\|Te_j\|=1$, $\forall 1\leq j\leq d$, then there exists a subset $\sigma \subseteq \{1, \dots, d\}$ of cardinality 
		\begin{align*}
			\operatorname{Card}(\sigma)\geq \frac{cd}{\|T\|^2}
		\end{align*}
		such that 
		\begin{align*}
			\left\|\sum_{j \in \sigma}a_jTe_j\right\|^2\geq A \sum_{j \in \sigma}|a_j|^2, \quad \forall a_j \in  \mathbb{K}, \forall j \in \sigma.
	\end{align*}}
\end{theorem}  
In 2009 Tropp gave a polynomial time algorithm for the proof Theorem \ref{BOURGAINTZARIRITHEOREM} \cite{TROPP}. It came as a surprise in 2009 (arXiv version) when  Spielman and Srivastava gave a simple proof of Theorem \ref{BOURGAINTZARIRITHEOREM} by proving a generalization of  Theorem \ref{BOURGAINTZARIRITHEOREM}  \cite{SPIELMANSRIVASTAVA} due to Vershynin \cite{VERSHYNIN}.  In 2012 Casazza and Pfander  proved infinite dimensional version of Theorem \ref{BOURGAINTZARIRITHEOREM} \cite{CASAZZAPFANDER}. For  beautiful descriptions of Bourgain-Tzafriri restricted invertibility theorem and generalizations we refer \cite{NAORYOUSSEF, CASAZZATREMAIN, NAOR, SRIVASTAVA, CASAZZA, YOUSSEF, RAVICHANDRAN, TROPP2, CHRETIENDARSES, YOUSSEF2, VERSHYNIN, MARCUSSPIELMANSRIVASTAVA, GLUSKINOLEVSKII}.

We formulate Conjectures \ref{RC} and \ref{NRC} which are  based on Theorem \ref{BOURGAINTZARIRITHEOREM}. Let $d\in \mathbb{N}$, $\mathcal{A}$ be a  unital C*-algebra and $\mathcal{A}^d$ be the left module over $\mathcal{A}$ w.r.t. natural operations. Modular $\mathcal{A}$-inner product on $\mathcal{A}^d$ is defined as 
  \begin{align*}
  	\langle (x_j)_{j=1}^d, (y_j)_{j=1}^d\rangle \coloneqq \sum_{j=1}^{d}x_jy_j^*,\quad \forall (x_j)_{j=1}^d, (y_j)_{j=1}^d \in \mathcal{A}^d.
  \end{align*}
  Hence the norm on $\mathcal{A}^d$ becomes 
  \begin{align*}
  	\|(x_j)_{j=1}^d\|\coloneqq \left\|\sum_{j=1}^{d}x_jx_j^*\right\|^\frac{1}{2}, \quad \forall (x_j)_{j=1}^d \in \mathcal{A}^d.
  \end{align*}
In this way $\mathcal{A}^d$ becomes \textbf{standard Hilbert C*-module} \cite{KAPLANSKY, PASCHKE, RIEFFEL}.  
  \begin{conjecture}\label{RC} 
  \textbf{[(Commutative) Modular Bourgain-Tzafriri Restricted Invertibility Conjecture]
  Let $\mathcal{A}$ be a  unital commutative C*-algebra and $\mathcal{I}(\mathcal{A})$ be the set of all invertible elements of $\mathcal{A}$.   For $d\in \mathbb{N}$, let $ \mathbb{M}_{d\times d}(\mathcal{A})$ be the set of all $d$ by $d$ matrices over $\mathcal{A}$. For  $ M\in \mathbb{M}_{d\times d}(\mathcal{A})$, let $\det (M) $ be the determinant of $M$. Let $\mathcal{A}^d$ be the standard (left) Hilbert C*-module over 	$\mathcal{A}$ and    $\{e_j\}_{j=1}^d$ be the  canonical orthonormal basis for $\mathcal{A}^d$. There are universal real  constants $A>0$, $c>0$ ($A$ and $c$ may  depend upon C*-algebra $\mathcal{A}$) satisfying the following property. If  $d \in \mathbb{N}$ and  $ M \in \mathbb{M}_{d\times d}(\mathcal{A})$    with  $\langle Me_j, Me_j \rangle =1$, $\forall 1\leq j\leq d$ and $\det (M) \in \mathcal{I}(\mathcal{A})\cup\{0\}$, then there exists a subset $\sigma \subseteq \{1, \dots, d\}$ of cardinality 
  \begin{align*}
  	\operatorname{Card}(\sigma)\geq \frac{cd}{\|M\|^2}
  \end{align*}
  such that 
  \begin{align*}
  	\sum_{j \in \sigma}\sum_{k \in \sigma}a_j \langle Me_j, Me_k\rangle a_k^*=\left\langle \sum_{j \in \sigma}a_jMe_j, \sum_{k \in \sigma}a_kMe_k\right\rangle \geq A \sum_{j \in \sigma}a_ja_j^*, \quad \forall a_j \in  \mathcal {A}, \forall j \in \sigma, 
  \end{align*}	
  where $\|M\|$ is the norm of the Hilbert C*-module homomorphism defined by $M$ as $M:\mathcal{A}^d \ni x \mapsto Mx \in \mathcal{A}^d$.}
   \end{conjecture}
  We next formulate Conjecture \ref{RC} for unital C*-algebras which need not be commutative. In the statement we use the notion of Manin matrices. We refer \cite{CHERVOVFALQUIRUBTSOV, CHERVOVFALQUI} for the basics of Manin matrices.
  \begin{conjecture}\label{NRC}
   \textbf{[(Noncommutative) Modular Bourgain-Tzafriri Restricted Invertibility Conjecture ]
  	Let $\mathcal{A}$ be a  unital  C*-algebra and $\mathcal{I}(\mathcal{A})$ be the set of all invertible elements of $\mathcal{A}$.   For $d\in \mathbb{N}$, let $ \mathbb{MM}_{d\times d}(\mathcal{A})$ be the set of all $d$ by $d$ Manin matrices over $\mathcal{A}$. For  $ M\in \mathbb{MM}_{d\times d}(\mathcal{A})$, let $\det^\text{column} (M) $ be the Manin determinant of $M$ by column expansion. Let $\mathcal{A}^d$ be the standard (left) Hilbert C*-module over 	$\mathcal{A}$ and    $\{e_j\}_{j=1}^d$ be the  canonical orthonormal basis for $\mathcal{A}^d$. There are universal real  constants $A>0$, $c>0$ ($A$ and $c$ may  depend upon C*-algebra $\mathcal{A}$) satisfying the following property. If  $d \in \mathbb{N}$ and  $ M \in \mathbb{MM}_{d\times d}(\mathcal{A})$    with  $\langle Me_j, Me_j \rangle =1$, $\forall 1\leq j\leq d$ and $\det^\text{column} (M) \in \mathcal{I}(\mathcal{A})\cup\{0\}$, then there exists a subset $\sigma \subseteq \{1, \dots, d\}$ of cardinality 
  	\begin{align*}
  		\operatorname{Card}(\sigma)\geq \frac{cd}{\|M\|^2}
  	\end{align*}
  	such that 
  	\begin{align*}
  		\sum_{j \in \sigma}\sum_{k \in \sigma}a_j \langle Me_j, Me_k\rangle a_k^*=\left\langle \sum_{j \in \sigma}a_jMe_j, \sum_{k \in \sigma}a_kMe_k\right\rangle \geq A \sum_{j \in \sigma}a_ja_j^*, \quad \forall a_j \in  \mathcal {A}, \forall j \in \sigma, 
  	\end{align*}	
  	where $\|M\|$ is the norm of the Hilbert C*-module homomorphism defined by $M$ as $M:\mathcal{A}^d \ni x \mapsto Mx \in \mathcal{A}^d$.} 
  \end{conjecture}
 \begin{remark}
 	\begin{enumerate}[\upshape (i)]
 		\item \textbf{We can surely formulate Conjecture \ref{RC} by removing the condition $\det (M) \in \mathcal{I}(\mathcal{A})\cup\{0\}$ and Conjecture \ref{NRC} by removing the condition Manin matrices and $\det^\text{column} (M) \in \mathcal{I}(\mathcal{A})\cup\{0\}$. But we strongly  believe that Conjectures  \ref{RC} and \ref{NRC}  fail with this much of generality.}
 		\item \textbf{If Conjecture \ref{RC} holds but Conjecture \ref{NRC} fails,  then we  can try Conjecture \ref{NRC} for W*-algebras or C*-algebras with invariant basis number (IBN) property.} We refer \cite{GIPSON} for the IBN properties of C*-algebras.
 	\end{enumerate}
 \end{remark}

  \section{Modular Johnson-Lindenstrauss Flattening Conjecture}
   Everything starts from the following surprising result of Prof. Johnson and Prof. Lindenstrauss from 1984 \cite{JOHNSONLINDENSTRAUSS}.
  \begin{theorem}\cite{MATOUSEKBOOK, JOHNSONLINDENSTRAUSS}\label{JL}
  	\textbf{(Johnson-Lindenstrauss Flattening Lemma - original form) Let $M, N\in \mathbb{N}$ and  $\mathbf{x}_1, \mathbf{x}_2, \dots, \mathbf{x}_M\in \mathbb{R}^N$. For each $0<\varepsilon<1$, there exists a Lipschitz map $f: \mathbb{R}^N\to \mathbb{R}^m$  and a real $r>0$ such that 
  		\begin{align*}
  			r(1-\varepsilon)\|\mathbf{x}_j-\mathbf{x}_k\|\leq \|f(\mathbf{x}_j)-f(\mathbf{x}_k)\|\leq r(1+\varepsilon)\|\mathbf{x}_j-\mathbf{x}_k\|, \quad \forall 1 \leq j, k\leq M, 
  		\end{align*}
  		where 
  		\begin{align*}
  			m=O\left(\frac{\log M}{\varepsilon^2}\right).
  	\end{align*}}
  \end{theorem}
Over the time, several improvements of Theorem \ref{JL} were obtained. We recall them. 
  \begin{theorem}\cite{FRANKLMAEHARA, JOHNSONLINDENSTRAUSS} \textbf{(Johnson-Lindenstrauss  Flattening  Lemma - Frankl-Maehara form)
  		Let $0<\varepsilon<\frac{1}{2}$   and $ M\in \mathbb{N}$. Define 
  		\begin{align*}
  			m (\varepsilon, M)\coloneqq \left\lceil 9\frac{1}{\varepsilon^2-\frac{2\varepsilon^3}{3}}\log M \right \rceil +1.
  		\end{align*} 
  		If  $M> m (\varepsilon, M)$, then for any  $\mathbf{x}_1, \mathbf{x}_2, \dots, \mathbf{x}_M\in \mathbb{R}^M$,  there exists 	a map $f: \{\mathbf{x}_j\}_{j=1}^M\to \mathbb{R}^m$   such that 
  		\begin{align*}
  			(1-\varepsilon)\|\mathbf{x}_j-\mathbf{x}_k\|^2\leq \|f(\mathbf{x}_j)-f(\mathbf{x}_k)\|^2\leq (1+\varepsilon)\|\mathbf{x}_j-\mathbf{x}_k\|^2, \quad \forall 1 \leq j, k\leq M.
  	\end{align*}}
  \end{theorem}
  \begin{theorem}\cite{DASGUPTAGUPTA, JOHNSONLINDENSTRAUSS}
  	\textbf{(Johnson-Lindenstrauss Flattening Lemma - Dasgupta-Gupta form)	
  		Let $M, N\in \mathbb{N}$ and  $\mathbf{x}_1, \mathbf{x}_2, \dots, \mathbf{x}_M\in \mathbb{R}^N$. Let $0<\varepsilon<1$. Choose any natural number $m$ such that 
  		\begin{align*}
  			m>4\frac{1}{\frac{\varepsilon^2}{2}-\frac{\varepsilon^3}{3}}\log M. 
  		\end{align*}
  		Then there exists a  map $f: \mathbb{R}^N\to \mathbb{R}^m$   such that 
  		\begin{align*}
  			(1-\varepsilon)\|\mathbf{x}_j-\mathbf{x}_k\|^2\leq \|f(\mathbf{x}_j)-f(\mathbf{x}_k)\|^2\leq (1+\varepsilon)\|\mathbf{x}_j-\mathbf{x}_k\|^2, \quad \forall 1 \leq j, k\leq M.
  		\end{align*}
  		The map $f$ can be found in randomized polynomial time.}
  \end{theorem}
  \begin{theorem} \cite{JOHNSONLINDENSTRAUSS, FOUCARTRAUHUT} \textbf{(Johnson-Lindenstrauss Flattening Lemma - matrix form)
  		There is a universal constant $C>0$ satisfying the following. Let $0<\varepsilon<1$, $M,N \in \mathbb{N}$ and $\mathbf{x}_1, \mathbf{x}_2, \dots, \mathbf{x}_M\in \mathbb{R}^N$. For each natural number 
  		\begin{align*}
  			m>\frac{C}{\varepsilon^2}\log M,
  		\end{align*}
  		there exists a matrix $M\in \mathbb{M}_{m\times N}(\mathbb{R})$ such that 
  		\begin{align*}
  			(1-\varepsilon)\|\mathbf{x}_j-\mathbf{x}_k\|^2\leq \|M(\mathbf{x}_j-\mathbf{x}_k)\|^2\leq (1+\varepsilon)\|\mathbf{x}_j-\mathbf{x}_k\|^2, \quad \forall 1 \leq j, k\leq M.
  	\end{align*}}
  \end{theorem}
  In 2016 Larsen and Nelson\cite{LARSENNELSON, LARSENNELSON2017} showed that bound in Johnson-Lindenstrauss flattening lemma is optimal. For a comprehensive look on Johnson-Lindenstrauss flattening lemma from the historical point of view and its various applications we refer important papers and books  \cite{ACHLIOPTAS, INDYKMOTWANI, KRAHMERWARD, INDYKNAOR, ALON, COHENELDERMUSCOMUSCOPERSU, BECCHETTIBURYCOHENADDADGRANDONISCHWIEGELSHOHN, JAYRAMWOODRUFF, OSTROVSKII, LINIALLONDONRABINOVICH, DEDEZALAURENT, BOURGAIN, JACQUES, INDYK2000, AILONCHAZELLE, AILONCHAZELLE2009, KANENELSON, DASGUPTAKUMARSARLOS, BARTALRECHTSCHULMAN, KANEMEKANELSON, KLARTAGMENDELSON, ARRIAGAVEMPALA, AILONLIBERTY, AILONLIBERTY2, ALONKLARTAG, FREKSEN, VENKATASUBRAMANIANWANG, LIBERTYAILONSINGER}. \\
  We now  formulate the following interesting problem. 
  \begin{problem}\label{MODULARPROBLEM}
  		\textbf{Let $\mathscr{A}$ be the set of all unital C*-algebras. What is the best function $\phi: \mathscr{A}\times (0,1)\times \mathbb{N} \to (0,\infty)$ satisfying the following. Let $\mathcal{A}$ be a unital C*-algebra. 	There is a universal constant $C>0$ (which may depend upon $\mathcal{A}$) satisfying the following. Let $0<\varepsilon<1$, $M,N \in \mathbb{N}$ and $\mathbf{x}_1, \mathbf{x}_2, \dots, \mathbf{x}_M\in \mathcal{A}^N$. For each natural number 
  	\begin{align*}
  		m>C\phi (\mathscr{A}, \varepsilon, M)
  	\end{align*}
  	there exists a matrix $M\in \mathbb{M}_{m\times N}(\mathcal{A})$ such that 
  	\begin{align*}
  		(1-\varepsilon)\langle \mathbf{x}_j-\mathbf{x}_k, \mathbf{x}_j-\mathbf{x}_k\rangle \leq \langle M(\mathbf{x}_j-\mathbf{x}_k), M(\mathbf{x}_j-\mathbf{x}_k)\rangle \leq (1-\varepsilon)\langle \mathbf{x}_j-\mathbf{x}_k, \mathbf{x}_j-\mathbf{x}_k\rangle, \quad \forall 1 \leq j, k\leq M.
  \end{align*}}
  \end{problem}
   A particular case of Problem \ref{MODULARPROBLEM} is the following  conjecture.
 \begin{conjecture}\label{MJLC}
	\textbf{(Modular Johnson-Lindenstrauss Flattening Conjecture)
		Let $\mathcal{A}$ be a unital C*-algebra. 	There is a universal constant $C>0$ (which may depend upon $\mathcal{A}$) satisfying the following. Let $0<\varepsilon<1$, $M,N \in \mathbb{N}$ and $\mathbf{x}_1, \mathbf{x}_2, \dots, \mathbf{x}_M\in \mathcal{A}^N$. For each natural number 
		\begin{align*}
			m>\frac{C}{\varepsilon^2}\log M,
		\end{align*}
		there exists a matrix $M\in \mathbb{M}_{m\times N}(\mathcal{A})$ such that 
		\begin{align*}
			(1-\varepsilon)\langle \mathbf{x}_j-\mathbf{x}_k, \mathbf{x}_j-\mathbf{x}_k\rangle \leq \langle M(\mathbf{x}_j-\mathbf{x}_k), M(\mathbf{x}_j-\mathbf{x}_k)\rangle \leq (1-\varepsilon)\langle \mathbf{x}_j-\mathbf{x}_k, \mathbf{x}_j-\mathbf{x}_k\rangle, \quad \forall 1 \leq j, k\leq M.
	\end{align*}}	
\end{conjecture}  
\begin{remark}
	\begin{enumerate}[\upshape(i)]
		\item We believe that Conjecture \ref{MJLC}  holds  at least for W*-algebras (von Neumann algebras) or C*-algebras with IBN property. 	
		\item 	\textbf{Modular Welch bounds} are derived in \cite{MAHESHKRISHNA} and various problems including modular Zauner's conjecture are stated there. 
	\end{enumerate}
\end{remark}

 \bibliographystyle{plain}
 \bibliography{reference.bib}

\end{document}